ID |
Date |
Author |
Target Material |
Target Oven W.O# |
Source |
Status |
Subject |
Text |
 |
189
|
Monday, May 04, 2015, 17:26 |
John Wong | | 40282-4 | TARGET BOAT | In progress | Sintering four different ZrC samples |
|
|
188
|
Friday, May 01, 2015, 10:50 |
Laura Lambert | | 40282-4 | TARGET BOAT | In progress | Sintering four different ZrC samples | Friday, May 01, 2015, 10:27
Four different samples of ZrC are
loaded into a target boat. |
 |
187
|
Thursday, April 23, 2015, 10:17 |
John Wong | SiC | 38817-7 | HP-SIS | Done | Sintering SiC#31 |
|
|
186
|
Friday, April 17, 2015, 16:14 |
Laura Lambert | SiC | 38817-7 | HP-SIS | In progress | Sintering SiC#31 | Friday, April 17, 2015, 16:01
Total number of foils: 520 discs |
|
185
|
Thursday, April 09, 2015, 09:38 |
Laura Lambert | Empty | 38817-7 | HP-SIS | Done | Sintering TaC coating container for SiC#32 | Thursday, April 09, 2015, 09:33
Target container is coated with |
|
184
|
Monday, March 16, 2015, 17:38 |
John Wong | Ta | 38817-5 | HP-SIS | Done | Conditioning Ta target |
|
|
183
|
Monday, March 09, 2015, 16:50 |
Laura Lambert | Ta | 38817-5 | HP-SIS | In progress | Conditioning Ta target | Monday, March 09, 2015, 16:47
Target consists of 400 discs.
Total packed length 3.4cm |
 |
182
|
Thursday, February 19, 2015, 15:21 |
Laura Lambert | Empty | 41177-2 | LP-SIS | Done | Sintering TaC coating container for UCx#12 | Thursday, February 19, 2015, 15:19
Target container coated with TaC,
loaded in EVAP2. |
|
181
|
Wednesday, February 11, 2015, 16:12 |
Laura Lambert | Empty | 41077-2 | LP-SIS | Done | Conditioning graphite insert inside container - with heat shield | Wednesday, February 11, 2015, 16:10
Container loaded into EVAP2.
3xTa heat shields. |
    |
180
|
Tuesday, February 10, 2015, 14:01 |
Laura Lambert | Empty | 41077-2 | LP-SIS | Done | Conditioning graphite insert inside container - no heat shield | Tuesday, February 10, 2015, 14:00
Container loaded into EVAP2 at noon.
no heat shields. |
  |
179
|
Tuesday, February 10, 2015, 13:19 |
Laura Lambert | Empty | 41077-2 | LP-SIS | Done | Coating/Sintering TaC on target container - For Graphite Insert |
|
|
178
|
Thursday, February 05, 2015, 18:03 |
John Wong | Empty | 41077-2 | LP-SIS | In progress | Coating/Sintering TaC on target container - For Graphite Insert |
|
|
177
|
Tuesday, February 03, 2015, 10:31 |
Laura Lambert | Empty | 41077-2 | LP-SIS | In progress | Coating/Sintering TaC on target container - For Graphite Insert |
|
|
176
|
Monday, February 02, 2015, 16:33 |
John Wong | Empty | 41077-2 | LP-SIS | In progress | Coating/Sintering TaC on target container - For Graphite Insert | Monday, February 02, 2015, 16:28
Container is coated with TaC.
No Ta foil wrapped around. |
|
175
|
Thursday, January 22, 2015, 14:21 |
Anders Mjos | Empty | | | Problems resolved | CG1A replaced | See previous e-log entry
for details |
|
174
|
Wednesday, January 14, 2015, 13:46 |
Anders Mjos | Empty | | | Maintenance | Pumping down for water leak test | Wednesday, January 14, 2015, 13:46:
Roughing
Wednesday, January 14, 2015, 14:57: |
|
173
|
Thursday, December 11, 2014, 10:20 |
Laura Lambert | | | | Maintenance | water supply shut off for lab closure over holidays | Thursday, December 11, 2014, 10:17
water supply and return valves closed
for lab closure over the holidays. |
|
172
|
Wednesday, November 26, 2014, 14:09 |
Laura Lambert | | | | Maintenance | OMEGA-IR dark patch and backing pump trip | Wednesday, November
26, 2014, 13:47 |
|
171
|
Wednesday, November 19, 2014, 15:10 |
Laura Lambert | Empty | 39530-2 | LP-FEBIAD | Done | test#1. direct voltage and OMEGA temperature measurements | Wednesday, November 19, 2014, 15:10
currently pumping down
Wednesday, November |
![test#1_TGHT_[without_vacuum].JPG test#1_TGHT_[without_vacuum].JPG](attachment.png)   |
170
|
Monday, November 10, 2014, 13:55 |
Laura Lambert | Empty | | LP-SIS | Done | Sintering TaC coating container for UCx#11 [2nd container] | Monday, November 10, 2014, 13:51
target container was sprayed with
TaC on friday and sintered over the weekend. |
|